OpenStudy (anonymous):

parallelogram create a equation to find for X insert X solution to find line AB and CD

OpenStudy (anonymous):

 also congruent or not congruent?

OpenStudy (anonymous):

doesnt congruent mean the same?

OpenStudy (campbell_st):

opposite sides of a parallelogram are equal so 5x + 2 = 8x - 7 now solve for x, when you get x you can find AB and CD by substituting and calculating... hope it helps

OpenStudy (anonymous):

x = 3

OpenStudy (anonymous):

that should help u more :)

OpenStudy (anonymous):

would my equation to solve for x then be 5(3)+2 = 8x-7?

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

no... yo solve you equation.. step 1 add 7 to both sides of the equation 5x + 2 = 8x - 7 +7 + 7 --------------- 5x + 9 = 8x now subtract 5 x from both sides of the equation 5x + 9 = 8x -5x -5x --------------- 9 = 3x now solve for x

OpenStudy (anonymous):

X=3 so then the equation above is the equation for x and this one 5(3)+2 = 8(3)-7 is to find AB and CD

OpenStudy (campbell_st):

thats correct... both sides have length 17 units